# A Generalized Alternating Anderson Acceleration Method

### v1

@@ 4.4.3 Nonnegative least squares problem
constrained with $\mathbf{x}\geq 0$. Here, the matrix $\mathbf{C}\in\mathbb{R}^{n_{1}\times n_{2}}$ and the vector $\widehat{\mathbf{x}}\in\mathbb{R}^{n_{1}}$ are given. The nonnegative least squares problem is an example of nonlinear problem with inequality constraints; see, e.g., [21].
@@
For this example, we set $n_{1}=150$, $n_{2}=300$, and take $\mathbf{C}$ as a randomly generated sparse matrix with density 0.01. Further, the vector $\widehat{\mathbf{x}}$ is randomly generated and sampled from the normal distribution. Finally, we set the penalty parameter $\mu=2$. For this test, we look for a reduction of $10^{-12}$ in the residual, allowing for a maximum of 2000 iterations. In Table 8, we report the number of iterations and the CPU time.
Table 8: Nonnegative least squares problem. Number of iterations and CPU time for ADMM, AA, and aAA–FP methods.
@@
Again, we note the beneficial effect of aAA(mm)[ss]–FP[$t$] over the simple ADMM. First, we note that AA($m$) is more than 10 times faster than ADMM, in terms of both number of iterations and CPU time, see, for example, AA(10). Furthermore, we observe that as $m$ increases, the number of iterations of both AA and aAA–FP decreases. However, no substantial gain is obtained for very large values of $m$; compare, for instance, AA(10) with AA($\infty$). We observe that, by carefully choosing the parameters $m,s,t$, one can construct an aAA(mm)[ss]–FP[$t$] solver that outperforms AA in terms of both number of iterations and CPU times. For example, aAA(10)[10]–FP[10] achieves the prescribed residual reduction in fewer iterations than AA(10), while requiring a CPU time comparable to that of AA(10) or AA($\infty$). Finally, as with the previous problems, we observe that aAA–FP can significantly accelerate the performance of ADMM, resulting in a solver that is more than 50 times faster than ADMM alone in terms of both number of iterations and CPU time; see, for example, aAA(5)[8]–FP[5] or aAA(8)[10]–FP[3].
@@ 4.5 Regularized logistic regression
where $\eta$ is the step length taken. We test how aAA(mm)[ss]–FP[$t$] can accelerate the convergence of GD for the (scaled) dataset covtype of the LIBSVM library [13] available at http://www.csie.ntu.edu.tw/~cjlin/libsvm; for this test set, we have $n_{1}=581,012$ and $n_{2}=54$. We would like to mention that a study of how AA($m$) and aAA(mm)[1]–FP[$m$] (referred to as AAP($m$) in [20]) accelerate GD for this problem was performed in [20]. The authors in [20] observed that the alternating Anderson–Picard presents a staircase pattern, being able to significantly reduce the error at each periodic AA; they observed that AAP(5) and AAP(7) converge to a tolerance of $10^{-10}$ in about 80 Picard iterations for the problem considered here, where they did not count the AA step in each period. Thus, AAP(5) and AAP(7) take around 96 and 91 iterations, respectively. We run GD, AA, and aAA–FP for 1000 iterations, seeking a reduction of $10^{-12}$ in the residual. As in [20], we set $\beta=10^{-2}$ and $\eta=1$. In Table 9, we report the number of iterations and the CPU time required to reach the prescribed tolerance.
Table 9: Regularized logistic regression. Number of iterations and CPU time for GD, AA, and aAA–FP methods.
@@ 5 Conclusion
In this work, we proposed a generalization of the alternating Anderson acceleration iteration, aAA(mm)[ss]–FP[$t$]. The proposed method consists of a periodic pattern of $t$ fixed point iterations followed by $s$ iterations of AA($m$). We provided a theoretical analysis of the convergence of the method, when applied for accelerating a linear fixed-point iteration. We established a connection between aAA($\infty$)[1]–FP[$t$] and GMRES in the linear case. Further, we gave a sufficient condition for the convergence of aAA–FP, when the fixed point iteration matrix is diagonalizable and noncontractive. Numerical results validated our theoretical findings, showing a periodic equivalence between GMRES and aAA($\infty$)[1]–FP[$t$]. In addition, we applied the proposed aAA–FP method to accelerate Jacobi iteration, Picard iteration, gradient descent, and the alternating direction method of multipliers for the solutions of problems arising in fluid dynamics and nonlinear optimization. The results demonstrated that aAA–FP can drastically accelerate the fixed-point iterations and achieve competitive performance with AA($\infty$), in terms of both number of iterations and CPU times, provided that $m$, $s$, and $t$ are appropriately selected.
From the numerical tests, we observed that the least squares problem to be solved within AA may be rank-deficient. Future work will focus on the development of efficient solvers for the least-squares problem that arises at each Anderson iteration. In addition, we mention that the code we employ in this work is not optimized. For this reason, future work will be devoted to devising Python and C++ implementations of the proposed strategy. Finally, we plan to explore the application of aAA–FP to more complex, real-world problems.
Acknowledgments

### v2

@@ 4.4.3 Nonnegative least squares problem
constrained with $\mathbf{x}\geq\mathbf{0}$. Here, the matrix $\mathbf{C}\in\mathbb{R}^{n_{1}\times n_{2}}$ and the vector $\widehat{\mathbf{x}}\in\mathbb{R}^{n_{1}}$ are given. The nonnegative least squares problem is an example of nonlinear problem with inequality constraints; see, e.g., [22].
@@
For this example, we set $n_{1}=150$, $n_{2}=300$, and take $\mathbf{C}$ as a randomly generated sparse matrix with density 0.01. Further, the vector $\widehat{\mathbf{x}}$ is randomly generated and sampled from the normal distribution. Finally, we set the penalty parameter $\mu=2$. For this test, we look for a reduction of $10^{-12}$ in the residual, allowing for a maximum of 2000 iterations.
In Table 9, we report number of iterations and CPU time for ADMM, AA, and aAA–FP methods using the zero vector as the initial guess. Then, we run the solvers for 50 randomly generated initial guesses that are generated by Matlab’s rand function with seed 100,000, and report the average number of iterations and the average CPU time in Table 10. Overall, the average CPU time is lower than that obtained with a single initial guess.
Table 9: Nonnegative least squares problem. Number of iterations and CPU time for ADMM, AA, and aAA–FP methods using the zero vector as the initial guess. The notation $\dagger$ indicates that the solver did not achieve the desired stopping criterion within 2000 iterations.
@@
Table 10: Nonnegative least squares problem. The average of number of iterations and the average CPU time for ADMM, AA, and aAA–FP methods using 50 random initial guesses generated by Matlab’s rand function with seed 100,000. The notation $\dagger$ indicates that the solver did not achieve the desired stopping criterion within 2000 iterations.
ADMM AA(1) AA(3) AA(10) AA($\infty$)
it CPU it CPU it CPU it CPU it CPU
[…]
it CPU it CPU it CPU it CPU it CPU
51 0.05 92 0.10 53 0.06 45 0.05 68 0.20
it CPU it CPU it CPU it CPU it CPU
46 0.05 61 0.07 55 0.06 43 0.06 46 0.06
it CPU it CPU it CPU it CPU it CPU
44 0.05 90 0.09 51 0.06 46 0.06 173 0.69
it CPU it CPU it CPU it CPU it CPU
43 0.06 93 0.10 52 0.06 43 0.06 47 0.06
In Table 9, we note the beneficial effect of aAA(mm)[ss]–FP[$t$] over the simple ADMM. First, we note that AA($m$) is more than 10 times faster than ADMM, in terms of both number of iterations and CPU time, see, for example, AA(10). Furthermore, we observe that as $m$ increases, the number of iterations of both AA and aAA–FP decreases. However, no substantial gain is obtained for very large values of $m$; compare, for instance, AA(10) with AA($\infty$). We observe that, by carefully choosing the parameters $m,s,t$, one can construct an aAA(mm)[ss]–FP[$t$] solver that outperforms AA in terms of both number of iterations and CPU times. For example, in Table 10, aAA(3)[8]–FP[8] achieves the prescribed residual reduction in fewer iterations than AA(3), while requiring a CPU time comparable to that of AA(10). Finally, as with the previous problems, we observe that aAA(mm)[ss]–FP[$t$] can significantly accelerate the performance of ADMM, resulting in a solver that is over 50 times faster than ADMM in iteration count and 34 times faster in CPU time; see, for example, aAA(5)[8]–FP[5] or aAA(8)[10]–FP[3] in Table 10.
In Figure 10 we plot the history of the residuals for aAA(10)[10]–FP(10), for the 50 different initial guesses. As observed, for some initial guesses the solver requires more than 50 iterations in order to reach convergence. This is not surprising, as the behavior of the classical Anderson acceleration strongly depends on the initial guess; see [18]. Nonetheless, the solver requires in average 43 iterations for reaching the prescribed reduction in the residual; see Table 10.
Figure 10: Nonnegative least squares problem. History of residuals for aAA(10)[10]–FP(10) for 50 random initial guesses generated by Matlab’s rand function with seed 100,000.
We would like to mention that for AA($\infty$) Matlab failed to solve the least-squares problem (4) for some of the instances because the matrix of the least-squares problem was found to be rank-deficient. In some cases, the first 15 columns (each column has the form $r(\mathbf{x}_{k})-r(\mathbf{x}_{k-i}$)) have very small norms, on the order of $10^{-9}$ or $10^{-10}$, which results in an effective loss of rank. In the literature, many efforts have been made to address how to efficiently and robustly solve the least-squares problems involved in AA, for example [28, 29]. We leave such implementation details for future work. In practice, we do not recommend using AA($\infty$). For aAA(mm)[ss]–FP[$t$], although the choice of $m$ is often problem-dependent, it should remain modest in size–a few tens–to ensure efficiency and stability.
@@ 4.5 Regularized logistic regression
where $\eta$ is the step length taken. We test how aAA(mm)[ss]–FP[$t$] can accelerate the convergence of GD for the (scaled) dataset covtype of the LIBSVM library [13] available at http://www.csie.ntu.edu.tw/~cjlin/libsvm; for this test set, we have $n_{1}=581,012$ and $n_{2}=54$. We would like to mention that a study of how AA($m$) and aAA(mm)[1]–FP[$m$] (referred to as AAP($m$) in [21]) accelerate GD for this problem was performed in [21]. The authors in [21] observed that the alternating Anderson–Picard presents a staircase pattern, being able to significantly reduce the error at each periodic AA; they observed that AAP(5) and AAP(7) converge to a tolerance of $10^{-10}$ in about 80 Picard iterations for the problem considered here, where they did not count the AA step in each period. Thus, AAP(5) and AAP(7) take around 96 and 91 iterations, respectively. We run GD, AA, and aAA–FP for 1000 iterations, seeking a reduction of $10^{-12}$ in the residual. As in [21], we set $\beta=10^{-2}$ and $\eta=1$. In Table 11, we report the number of iterations and the CPU time required to reach the prescribed tolerance. For this test, we employ the zero vector as the initial guess.
Table 11: Regularized logistic regression. Number of iterations and CPU time for GD, AA, and aAA–FP methods, when employing the zero vector as the initial guess. The notation $\dagger$ indicates that the solver did not achieve the desired stopping criterion within 1000 iterations.
@@
4.6 Study the parameters choice
We study now the choice of the parameters, $m,s$ and $t$ in aAA(mm)[ss]–FP[$t$]. We consider the total variation problem described in Section 4.4.1, and run aAA(mm)[ss]–FP[$t$], employing zero as the initial guess. We report the number of iterations and the CPU times in second for three different cases:
  1. (1.)
    set $m=s$ and let $m,t\in\{1,\ldots,20\}$;
  2. (2.)
    set $m=t$ and let $m,s\in\{1,\ldots,20\}$;
  3. (3.)
    set $s=t$ and let $m,t\in\{1,\ldots,20\}$.
In Figures 11, 12, and 13 we report the contour plots of the number of iterations and the CPU times. In [4], it was found that taking $m\geq 2t$ generally improves performance of aAA(mm)[1]–FP[$t$] for the problems considered. Accordingly, we have also included the line $m=2t$ in the contour plots. First, from Figure 11 we observed that in both terms of total number of aAA(mm)[$m$]-FP[$t$] iterations and CPU time it is more beneficial increasing the number of Anderson iterations $s$ ($m=s$) than increasing the number of fixed point iterations $t$. Second, Figure 12 shows that increasing the number of Anderson iterations ($s$) results in a lower number of aAA(mm)[$s$]-FP[$m$] iterations and computational time than increasing the number of fixed point iterations ($t=m$). Third, as observed in Figure 13 one can obtain a reduction in the number of aAA(mm)[$t$]-FP[$t$] iterations and in the computational time by letting the window size $m$ increase. We remark the best choice of the parameters are inside the dark blue area. Figures 11 and 13 seem suggest that using $m\geq 2t$ gives a better performance.
The idea behind aAA(mm)[ss]–FP[$t$] is to apply $t$ fixed‑point iterations between every $s$ AA($m$) updates in order to reduce the computational cost of the least‑squares problems. However, if $t$ is too large, the method behaves more like a standard fixed‑point iteration and the convergence may slow down. Thus, there is a trade‑off between iteration count and CPU time. In general, we recommend that $t$ not be chosen significantly larger than the parameters $m$ and $s$, since these have a more substantial impact on performance.
Figure 11: Total variation. Contour plot of the number of iterations (left) and of the CPU times (right), when applying aAA(mm)[$m$]-FP[$t$], for $m,t\in\{1,\ldots,20\}$. The red line corresponds to $m=2t$.
Figure 12: Total variation. Contour plot of the number of iterations (left) and of the CPU times (right), when applying aAA(mm)[$s$]-FP[$m$], for $m,s\in\{1,\ldots,20\}$.
Figure 13: Total variation. Contour plot of the number of iterations (left) and of the CPU times (right), when applying aAA(mm)[$t$]-FP[$t$], for $m,t\in\{1,\ldots,20\}$. The red line corresponds to $m=2t$.
@@ 5 Conclusion
In this work, we proposed a generalization of the alternating Anderson acceleration iteration, aAA(mm)[ss]–FP[$t$]. The proposed method consists of a periodic pattern of $t$ fixed point iterations followed by $s$ iterations of AA($m$). We provided a theoretical analysis of the convergence of the method, when applied for accelerating a linear fixed-point iteration. We established a connection between aAA($\infty$)[1]–FP[$t$] and GMRES in the linear case. Further, we gave a sufficient condition for the convergence of aAA–FP, when the fixed point iteration matrix is diagonalizable and noncontractive. Numerical results validated our theoretical findings, showing a periodic equivalence between GMRES and aAA($\infty$)[1]–FP[$t$]. In addition, we applied the proposed aAA–FP method to accelerate Jacobi iteration, Gauss–Seidel iteration, Picard iteration, gradient descent, and the alternating direction method of multipliers for the solutions of problems arising in fluid dynamics and nonlinear optimization. The results demonstrated that aAA–FP can drastically accelerate the fixed-point iterations and achieve competitive performance with AA($\infty$), in terms of both number of iterations and CPU times, provided that $m$, $s$, and $t$ are appropriately selected.
From the numerical tests, we observed that the least squares problem to be solved within AA may be rank-deficient for large window size $m$. Future work will focus on the development of efficient solvers for the least-squares problem that arises at each Anderson iteration. In addition, we mention that the code we employ in this work is not optimized. For this reason, future work will be devoted to devising Python and C++ implementations of the proposed strategy. Finally, we plan to explore the application of aAA–FP to more complex, real-world problems.
Acknowledgements